A crane with an arm length of 20 meters moves through an angle of 25 degrees while lifting a load. What is the angle it moved through in radians?
1
Expert's answer
2020-03-17T09:52:48-0400
"\theta=25\frac{2\pi}{360}=\frac{5\pi}{36}\ rad=0.436\ rad"